OpenStudy (anonymous):

A right triangle has one side of length 8 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

Pythagorean Theorem a^2 + b^2 = c^2 Do you know which one has to be c?

OpenStudy (anonymous):

No I do not

OpenStudy (anonymous):

the hypotneuse --------> 10 so c=10 and a = 8 substitute 8^2 + b^2 = 10^2 or 64 + b^2 = 100 do you know the next step?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

subtract 64 from both sides b^2 = 36 Next step?

OpenStudy (anonymous):

Do I simplify after this?

OpenStudy (anonymous):

Yes we have been simplifying ...

OpenStudy (anonymous):

So would you get b^2=36?

OpenStudy (anonymous):

and simplify this ...

OpenStudy (anonymous):

Would it be the square root of 36?

OpenStudy (anonymous):

yes. Nice job!

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Your welcome